# LazyReview A Dataset for Uncovering Lazy Thinking in NLP Peer Reviews

### v1

@@ Abstract
Peer review is a cornerstone of quality control in scientific publishing. With the increasing workload, the unintended use of ‘quick’ heuristics, referred to as lazy thinking, has emerged as a recurring issue compromising review quality. Automated methods to detect such heuristics can help improve the peer-reviewing process. However, there is limited NLP research on this issue, and no real-world dataset exists to support the development of detection tools. This work introduces LazyReview, a dataset of peer-review sentences annotated with fine-grained lazy thinking categories. Our analysis reveals that Large Language Models (LLMs) struggle to detect these instances in a zero-shot setting. However, instruction-based fine-tuning on our dataset significantly boosts performance by 10-20 performance points, highlighting the importance of high-quality training data. Furthermore, a controlled experiment demonstrates that reviews revised with lazy thinking feedback are more comprehensive and actionable than those written without such feedback. We will release our dataset and the enhanced guidelines that can be used to train junior reviewers in the community. [^1] Code available here: https://github.com/UKPLab/arxiv2025-lazy-review
@@ 2.2 Review Formatting
We utilize GPT-4 OpenAI et al. (2024) to prepare the instances (review segments) for subsequent annotations. We instruct GPT-4 to extract the review segments from the ‘Summary of Weaknesses’ section of the review that can be classified into one of the lazy thinking classes, as outlined in the ARR guidelines [^4] The prompt for GPT-4 is provided in Appendix §A.4. Since lazy thinking often contributes to reasons for rejecting a paper, we believe it will appear only in the ‘Summary of Weaknesses’ section. We obtain 1,776 review segments after this step, with each segment having varied lengths from 1 to 5 review sentences, as described later in Sec 2.5. To validate the quality of the extracted segments, we sample a set of $100$ segments from this pool and task three annotators to independently annotate each segment within the context of the entire review to decide on their candidacy towards lazy thinking. The final label is chosen based on a majority vote. We obtain Cohen’s $\kappa$ of 0.82 and obtain precision, recall and F1 scores as 0.74, 1.00 and 0.85, respectively. Intuitively, this means that GPT-4 samples more candidates than required, giving us a larger pool of candidates for annotation. We therefore introduce another class to our labelling schema, ‘None’ to annotate non-lazy thinking candidates.
@@ 2.4 Evolving Guidelines
Final Validation. To validate the guidelines developed over three rounds, we hired a new group of English-speaking PhD students with NLP peer review expertise to annotate the same set of instances. These annotators were not involved in the development of the guidelines. Re-annotation across rounds 1, 2, and 3 resulted in inter-annotator agreement ($\kappa$) values of 0.32, 0.36, and 0.48, respectively. This steady increase in $\kappa$ aligns with previous results of 0.31, 0.38, and 0.52, validating the iterative improvement of the guidelines based on stronger inter-annotator agreement. Agreement scores of 0.4 or higher are considered substantial due to the inherent subjectivity of the task and are consistent with agreement levels reported in previous studies within the peer-reviewing domain Kennard et al. (2022); Purkayastha et al. (2023). We retain the annotators from the initial rounds who developed the guidelines to annotate a total of 500 review segments as a part of our dataset, LazyReview. The overall annotation time amounted to 20 hours. This corresponds to 25 minutes per example, ranging from less than a minute for annotating shorter segments up to half an hour for longer ones.
@@ 3 Experiments
Tasks We propose two formulations for detecting lazy thinking in peer reviews: (i) Coarse-grained classification: a binary task to determine if an input segment $x$ is lazy thinking, and (ii) Fine-grained classification: a multi-class task to classify $x$ into one of the specific lazy thinking classes, $c_{i}\in C$.
Models. Since the guidelines for NLP conferences are mainly instructions, we explore various open-source instruction-tuned LLMs for this task. We experiment with the chat versions of LLaMa-2 7B and 13B Touvron et al. (2023) (abbr. LLaMa, LLaMaL), Mistral 7B instruct Jiang et al. (2023) (abbr. Mistral), Qwen-Chat 1.5 7B Bai et al. (2023) (abbr. Qwen), Yi-1.5 6B instruct AI et al. (2024) (abbr. Yi-1.5), Gemma-1.1 7B instruct Team et al. (2024) (abbr. Gemma), and SciTülu 7B Wadden et al. (2024) (abbr. SciTülu). [^6] The justification for the choice of these models along with implementation details are in Appendix §A.2
@@
We show the results using 1 static example in Table 3. ICL-based methods surpass zero-shot models, with notable gains in coarse classification (20 points for Gemma, 21 points for SciTülu string-based accuracy (S.A.)). We also obtain positive increments across the board in fine-grained classifications. SciTülu excels in the coarse-grained classification, possibly due to the domain specific pre-training. Qwen leads in fine-grained classification, likely due to the multi-lingual pre-training data and high-quality data filtering performed for the pre-training phase Bai et al. (2023).
@@
Table 4: Performance of LLMs after instruction tuning ($it$) for fine-grained classification using target segment (T) and the combination of review and target segment (RT) in terms of string-matching accuracy (St. (Acc)). The first row of each model states the best results obtained previously as detailed in Tables 2 and 3 respectively. Subscripts represent increment or decrement compared to the non-instruction tuned versions.
Results. We compare instruction-tuned models to their zero and few-shot counterparts for fine-grained and coarse-grained classification using 3-fold cross-validation, as shown in Tables 12 and 13 (cf. §A.10). Instruction tuning significantly enhances model performance. The LLaMa models and SciTülu excel with the SciRIFF Mix, benefiting from pre-training on SciRIFF and Tülu. Gemma and Qwen achieve their best results with the Tülu Mix, with Qwen outperforming Gemma (45.5 vs. 31.6 S.A. accuracy), likely due to Qwen’s multilingual training on 2.4T tokens Bai et al. (2023) compared to Gemma’s English-only training Team et al. (2024). Mistral and Yi perform best with the full dataset mix, with Yi surpassing Mistral (35.7 vs. 26.8 S.A.), possibly due to its larger vocabulary size. Qwen and Yi lead in fine-grained classification, while SciTülu excels in coarse classification, consistent with earlier findings. Instruction tuning’s effectiveness relies on data composition; blending general-purpose instruction data from Tülu with scientific data from SciRiFF optimizes LLM performance Shi et al. (2023) on this task. However, including tasks from all sources (Full Mix) can occassionally underperform, likely due to negative task transfer Jang et al. (2023); Wadden et al. (2024).
@@ 4 Related Work
Review Quality Analysis. As research outputs continue to rise, interest in assessing review quality has grown significantly within the academic community Kuznetsov et al. (2024). Previous studies have examined various aspects of review reports, including the importance of politeness and professional etiquette Bharti et al. (2023, 2024), thoroughness Severin et al. ; Guo et al. (2023), and comprehensiveness Yuan et al. (2022). There have been efforts to automate the peer-reviewing process using large language models Du et al. (2024); Zhou et al. (2024). However, there has been no focus on analyzing the usage of heuristics within the reviews. This work focuses on lazy thinking, a common issue affecting the quality of peer reviews in Natural Language Processing (NLP). While Rogers and Augenstein (2020) have qualitatively analyzed heuristics in NLP conferences, we are the first to present a dedicated dataset for identifying such practices and to propose automated methods for detecting low-quality reviews, thereby enhancing the effectiveness of the peer-review process in NLP.
@@ 5 Conclusion
In this work, we introduce LazyReview, a novel resource for detecting lazy thinking in NLP peer reviews. While the resource is aimed at promoting sound reviewing practices and enhancing overall review quality, it has some limitations. First, we adopt the definition and categories of lazy thinking from the ARR guidelines Rogers and Augenstein (2021) and the EMNLP Blog Liu et al. (2020), which are specific to NLP conference reviews. Therefore, the resource does not encompass peer-reviewing venues beyond ARR, and it would need adaptation to suit other domains with differing definitions of lazy thinking. Second, from a task and modeling perspective, we focus on the under-explored area of detecting lazy thinking in peer reviews. Although we conduct thorough experiments and analyses, the findings may not be fully generalizable to other classification tasks in the scientific domain, so results should be interpreted with caution. Additionally, ACL ARR guidelines currently characterize lazy thinking specifically within the weakness section of a review; future research could explore how such patterns manifest in other parts of the review as well. In addition, lazy thinking patterns may extend into subsequent author–reviewer discussions. However, as our starting dataset, NLPeer, does not include these interactions, we refrain from exploring this aspect, leaving it as a potential direction for future investigation. Furthermore, this study focuses exclusively on reviews authored prior to the widespread adoption of large language models (i.e., before 2023). As AI-generated reviews become increasingly prevalent, exploring their impact within the lazy thinking framework presents a compelling avenue for future research.
@@ Acknowledgements
This work has been funded by the German Research Foundation (DFG) as part of the Research Training Group KRITIS No. GRK $2$, along with the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts, within their joint support of the National Research Center for Applied Cybersecurity ATHENE. We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research). The work of Anne Lauscher is funded under the Excellence Strategy of the German Federal Government and the Federal States.
@@ A.2 Model and Implementation details
We select the LLMs based on multiple criteria: (a) all the models should be open-sourced in order to deploy them in real-world reviewing systems, (b) their size should be reasonable in order to perform full fine-tuning using LoRA Hu et al. (2022). c) they must be state-of-the-art on the existing LLM leaderboards Chiang et al. (2024). Based on these criteria, we select the chat version of the various models shown in Table 6. We use vLLM for fast inference of all the models. [^14] https://github.com/vllm-project/vllm. We set the temperature as 0 to have consistent predictions throughout and limit the output tokens to 30.
@@ A.3 Computational Budget
Gemma 1.1 Team et al. (2024) 7B google/gemma-1.1-7b-it
@@ A.9 Confusion matrices for Annotation Labels and Confidences across annotation rounds
We show the confusion matrices for annotation labels and confidences in Figures 16 and 17 respectively.
@@ A.11 Results for all the experiments across the annotation rounds
A.12 Instructions to Annotators for Rewriting Reviews and Evaluation
@@
Instruction for evaluating the rewrites You are provided a paper along with a pair of reviews written for the same paper. Your task is to perform pair-wise comparison of the reviews based on: (1) Constructiveness- reviewers should provide actionable feedback for the authors to improve on the paper, (2) Justified - reviewers should clearly state the reason for their arguments rather than putting out vague statements. Additionally, we introduce a measure, (3)A dherence, which judges how well the reviews are written based on the current ARR guidelines.

### v3

@@ Abstract
Peer review is a cornerstone of quality control in scientific publishing. With the increasing workload, the unintended use of ‘quick’ heuristics, referred to as lazy thinking, has emerged as a recurring issue compromising review quality. Automated methods to detect such heuristics can help improve the peer-reviewing process. However, there is limited NLP research on this issue, and no real-world dataset exists to support the development of detection tools. This work introduces LazyReview, a dataset of peer-review sentences annotated with fine-grained lazy thinking categories. Our analysis reveals that Large Language Models (LLMs) struggle to detect these instances in a zero-shot setting. However, instruction-based fine-tuning on our dataset significantly boosts performance by 10-20 performance points, highlighting the importance of high-quality training data. Furthermore, a controlled experiment demonstrates that reviews revised with lazy thinking feedback are more comprehensive and actionable than those written without such feedback. We will release our dataset and the enhanced guidelines that can be used to train junior reviewers in the community. [^1] Code available here: https://github.com/UKPLab/acl2025-lazy-review
@@ 2.2 Review Formatting
We utilize GPT-4 Achiam et al. (2024) to prepare the instances (review segments) for subsequent annotations. We instruct GPT-4 to extract the review segments from the ‘Summary of Weaknesses’ section of the review that can be classified into one of the lazy thinking classes, as outlined in the ARR guidelines [^4] The prompt for GPT-4 is provided in Appendix §A.4. Because lazy thinking is often a factor in paper rejection and is specifically defined within the ‘Summary of Weaknesses’ section according to the ARR guidelines Rogers and Augenstein (2021), we limit our focus to this section. Additionally, ARR serves as the primary reviewing platform for major NLP conferences, such as ACL, EMNLP, and NAACL. As a result, the reviews encompass a wide range of topics and discussions across various prominent NLP conferences. We obtain 1,776 review segments after this step, with each segment having varied lengths from 1 to 5 review sentences, as described later in Sec 2.5. To validate the quality of the extracted segments, we sample a set of $100$ segments from this pool and task three annotators to independently annotate each segment within the context of the entire review to decide on their candidacy towards lazy thinking. The final label is chosen based on a majority vote. We obtain Cohen’s $\kappa$ of 0.82 and obtain precision, recall and F1 scores as 0.74, 1.00 and 0.85, respectively. Intuitively, this means that GPT-4 samples more candidates than required, giving us a larger pool of candidates for annotation. We therefore introduce another class to our labelling schema, ‘None’ to annotate non-lazy thinking candidates.
@@ 2.4 Evolving Guidelines
Final Validation. To validate the guidelines developed over three rounds, we hired a new group of English-speaking PhD students with NLP peer review expertise to annotate the same set of instances. These annotators were not involved in the development of the guidelines. Re-annotation across rounds 1, 2, and 3 resulted in inter-annotator agreement ($\kappa$) values of 0.32, 0.36, and 0.48, respectively. This steady increase in $\kappa$ aligns with previous results of 0.31, 0.38, and 0.52, validating the iterative improvement of the guidelines based on stronger inter-annotator agreement. Agreement scores higher than 0.4 are considered moderate due to the inherent subjectivity of the task and are consistent with agreement levels reported in previous studies within the peer-reviewing domain Kennard et al. (2022); Purkayastha et al. (2023). We retain the annotators from the initial rounds who developed the guidelines to annotate a total of 500 review segments as a part of our dataset, LazyReview. The overall annotation time amounted to 20 hours. This corresponds to 25 minutes per example, ranging from less than a minute for annotating shorter segments up to half an hour for longer ones.
@@ 3 Experiments
Tasks. We propose two formulations for detecting lazy thinking in peer reviews: (i) Coarse-grained classification: a binary task to determine if an input segment $x$ is lazy thinking, and (ii) Fine-grained classification: a multi-class task to classify $x$ into one of the specific lazy thinking classes, $c_{i}\in C$.
Models. Since the guidelines for NLP conferences are mainly instructions, we explore various open-source instruction-tuned LLMs for this task. We experiment with the chat versions of LLaMa-2 7B and 13B Touvron et al. (2023) (abbr. LLaMa, LLaMaL), Mistral 7B instruct Jiang et al. (2023) (abbr. Mistral), Qwen-Chat 1.5 7B Bai et al. (2023) (abbr. Qwen), Yi-1.5 6B instruct Young et al. (2024) (abbr. Yi-1.5), Gemma-1.1 7B instruct Mesnard et al. (2024) (abbr. Gemma), and SciTülu 7B Wadden et al. (2024) (abbr. SciTülu). [^6] The justification for the choice of these models, along with implementation details, is in Appendix §A.2
@@
We show the results using 1 static example in Table 3. ICL-based methods surpass zero-shot models, with notable gains in coarse classification (20 points for Gemma, 21 points for SciTülu string-based accuracy (S.A.)). We also obtain positive increments across the board in fine-grained classifications. SciTülu excels in the coarse-grained classification, possibly due to the domain specific pre-training. Qwen leads in fine-grained classification, likely due to the multi-lingual pre-training data and high-quality data filtering performed for the pre-training phase Bai et al. (2023). [^11] Further analysis on the performance of Qwen is in Appendix A.15.
@@
Table 4: Performance of LLMs after instruction tuning ($it$) for fine-grained classification using target segment (T) and the combination of review and target segment (RT) in terms of string-matching accuracy (St. (Acc)). The first row of each model states the best results obtained previously as detailed in Tables 2 and 3 respectively. Subscripts represent increment or decrement compared to the non-instruction tuned versions. Increments compared to the first row of each model are highlighted in red and decrements or no change in gray.
Results. We compare instruction-tuned models to their zero and few-shot counterparts for fine-grained and coarse-grained classification using 3-fold cross-validation, as shown in Tables 12 and 13 (cf. §A.10). Instruction tuning significantly enhances model performance. The LLaMa models and SciTülu excel with the SciRIFF Mix, benefiting from pre-training on SciRIFF and Tülu. Gemma and Qwen achieve their best results with the Tülu Mix, with Qwen outperforming Gemma (45.5 vs. 31.6 S.A. accuracy), likely due to Qwen’s multilingual training on 2.4T tokens Bai et al. (2023) compared to Gemma’s English-only training Mesnard et al. (2024). Mistral and Yi perform best with the full dataset mix, with Yi surpassing Mistral (35.7 vs. 26.8 S.A.), possibly due to its larger vocabulary size. Qwen and Yi lead in fine-grained classification, while SciTülu excels in coarse classification, consistent with earlier findings. Instruction tuning’s effectiveness relies on data composition; blending general-purpose instruction data from Tülu with scientific data from SciRiFF optimizes LLM performance Shi et al. (2023) on this task. However, including tasks from all sources (Full Mix) can occassionally underperform, likely due to negative task transfer Jang et al. (2023); Wadden et al. (2024).
@@ 4 Related Work
Review Quality Analysis. As the number of publications continues to steeply grow, interest in assessing review quality has grown significantly within the academic community Kuznetsov et al. (2024). Previous studies have examined various aspects of review reports, including the importance of politeness and professional etiquette Bharti et al. (2023, 2024), thoroughness Severin et al. (2023); Guo et al. (2023), and comprehensiveness Yuan et al. (2022). There have been efforts to automate the peer-reviewing process using large language models Du et al. (2024); Zhou et al. (2024). However, there has been no focus on analyzing the usage of heuristics within the reviews. This work focuses on lazy thinking, a common issue affecting the quality of peer reviews in Natural Language Processing (NLP). While Rogers and Augenstein (2020) have qualitatively analyzed heuristics in NLP conferences, we are the first to present a dedicated dataset for identifying such practices and to propose automated methods for detecting low-quality reviews, thereby enhancing the effectiveness of the peer-review process in NLP.
@@ 5 Conclusion
In this work, we introduce LazyReview, a novel resource for detecting lazy thinking in NLP peer reviews. While the resource is aimed at promoting sound reviewing practices and enhancing overall review quality, it has some limitations. First, we adopt the definition and categories of lazy thinking from the ARR guidelines Rogers and Augenstein (2021) and the EMNLP Blog Liu et al. (2020), which are specific to NLP conference reviews. Therefore, the resource does not encompass peer-reviewing venues beyond ARR, and it would need adaptation to suit other domains and conferences such as ICLR with differing definitions of lazy thinking. Second, from a task and modeling perspective, we focus on the under-explored area of detecting lazy thinking in peer reviews. Although we conduct thorough experiments and analyses, the findings may not be fully generalizable to other classification tasks in the scientific domain, so results should be interpreted with caution. Additionally, ACL ARR guidelines currently characterize lazy thinking specifically within the weakness section of a review; future research could explore how such patterns manifest themselves in other parts of the review as well. In addition, lazy thinking patterns may extend into subsequent author–reviewer discussions. However, as our starting dataset, NLPeer, does not include these interactions, we refrain from exploring this aspect, leaving it as a potential direction for future investigation. Furthermore, this study focuses exclusively on reviews authored prior to the widespread adoption of large language models (i.e., before 2023). We hypothesize that the lazy thinking framework can be helpful in discriminating between human-written and AI-generated reviews.
@@ Acknowledgements
This work has been funded by the German Research Foundation (DFG) as part of the Research Training Group KRITIS No. GRK $2$. We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research). The work of Anne Lauscher is funded under the Excellence Strategy of the German Federal Government and the Federal States.
@@ A.2 Model and Implementation details
We select the LLMs based on multiple criteria: (a) all the models should be open-sourced and privacy-preserving in order to deploy them in real-world reviewing systems, (b) their size should be reasonable in order to perform full fine-tuning using LoRA Hu et al. (2022). c) they must be state-of-the-art on the existing LLM leaderboards Chiang et al. (2024). Based on these criteria, we select the chat version of the various models shown in Table 6. We use vLLM for fast inference of all the models. [^15] https://github.com/vllm-project/vllm. We set the temperature as 0 to have consistent predictions throughout and limit the output tokens to 30.
@@ A.3 Computational Budget
Gemma 1.1 Mesnard et al. (2024) 7B google/gemma-1.1-7b-it
@@ A.9 Confusion matrices for Annotation Labels and Confidences across annotation rounds
We show the confusion matrices for annotation labels and confidences in Figures 17 and 18 respectively.
@@ A.11 Results for all the experiments across the annotation rounds
A.12 Instructions to Annotators for Rewriting Reviews and Evaluation Protocol
@@
Instruction for evaluating the rewrites. You are provided a paper along with a pair of reviews written for the same paper. Your task is to perform pair-wise comparison of the reviews based on: (1) Constructiveness- reviewers should provide actionable feedback for the authors to improve on the paper, (2) Justified - reviewers should clearly state the reason for their arguments rather than putting out vague statements. Additionally, we introduce a measure, (3)Adherence, which judges how well the reviews are written based on the current ARR guidelines.
@@ A.14 Analysis on the Silver Annotations
A.15 Analysis on performance of model for different classes
We analyze the performance of the best-performing model, Qwen, on the Round 3 annotations of our dataset. Since this portion yielded the highest inter-annotator agreement, we specifically focus on how the model’s predictions distribute across classes here. The confusion matrix is shown in Fig 16. The model particularly struggles with the classes ‘Contradict expectations’ and ‘Negative results’, which mirror the difficulties observed in the human annotations from Round 3 (Fig 17(c)). On the other hand, ‘Language errors’ and ‘Niche topics’ remain among the easiest classes for the automated methods to identify.
Figure 16: Performance analysis of Qwen on Round 3 annotation data